2. Given the following demand for a taxi

Qdt = 50 – 0.4 +0.5pb-0.25Y

Pt = 20, Pb = 30, Y = 52

Where;

Qdt = Quantity demanded of a taxi

Pt = transport fare for a taxi

Pb = Transport fare for a bus

i) Calculate the price elasticity of demand for a taxi (3marks)

ii) Calculate the cross elasticity of demand between the two modes of transport (5marks)

iii) Calculate the income elasticity of demand (5marks);



1
Expert's answer
2021-04-11T20:05:07-0400
"Q_{dt}=50-0.4p_t+0.5p_b-0.25Y"

"Q_{dt}=44"

"E_{dt}=-0.4\times \frac {20}{44}=-0.182"

"E_{tb}=-\frac {0.4}{0.5}"

"E_Y=-0.25\times \frac{52}{44}=-0.295"